OpenStudy (anonymous):

A man 6 feet tall walks at a rate of 5 ft/sec toward a light that is 20 ft. above the ground. When he is 10 feet from the base of the ligh, at what rate is the tip of his shadow moving?

OpenStudy (amistre64):



OpenStudy (amistre64):

similar triangles gives us: 20/d+s = 6/s 20s = 6(d+s) this is the relation between the stuff i beleive

OpenStudy (amistre64):

[20s = 6(d+s)]' 20 s' = 6 (d'+s') 10s' = 3d'+3s' 10s'-3s' = 3d' s'(7) = 3d' s' = 3d'/7 is what I get for the rate of the shadow itself

OpenStudy (amistre64):

d' = 5 since the man is walking at a rate of 5 ....

OpenStudy (amistre64):

so if i did it right :) the shadow with respect to the mans walking is moving at 15/7 units per sec; and if we add that to the rate at which the person is walking we get: 15/7 + 5 ; and it might need to be negative since its towards the light

OpenStudy (amistre64):

id prolly feel better if we did this with trig instead of similar triangles; i always feel like im missing something.

OpenStudy (amistre64):

w = walking distance, and s = shadows length tan(a) = (w+s)/20 [20 tan(a) = w+s]' 20 sec^2(a) a' = w' + s' 20 sec^2(a) a' - w' = s' ewww, maybe not, since we dont know the rate of change of the angle, we could prolly find it out form the info, but it looks messier

OpenStudy (amistre64):

the concern i have is that this seems to work out the same no matter how far from the light he is since there is no place to plug in a value for the distance he is from the light ....

OpenStudy (amistre64):

\[\frac{ds}{dt}=\frac{ds}{dw}*\frac{dw}{dt}\] \[\frac{ds}{dt}=\frac{ds}{dw}*5\] as long as the rate of change of the shadow with repect to walking is 3/7 regardless of distance from the light, its good ; but i doubt that

OpenStudy (amistre64):

Dw [20s = 6(w+s)] 20s' = 6(1+s') 20s' = 6+6s' 20s'-6s' = 6 s'(14) = 6 = 3/7 ....

OpenStudy (anonymous):

Thanks a lot!